OpenStudy (anonymous):

A 12.82 g sample of a compound contains 4.09 g potassium (K), 3.71 g chlorine (Cl), and oxygen (O). Calculate the empirical formula.

OpenStudy (aaronq):

convert all the masses to moles, divide them all by the species with least moles. Normalize the moles into integers.

OpenStudy (anonymous):

So would I go 12.8*1/4.09 to find K?

OpenStudy (aaronq):

you have to convert the masses to moles first

OpenStudy (aaronq):

\(n=\dfrac{4.09 g}{Molar\;mass\;of\;K}\)

OpenStudy (anonymous):

So K is 39/1 mol so 4.09/39?

OpenStudy (aaronq):

yes, do that for Cl and for O, as well.

OpenStudy (anonymous):

.10487179 K, .10465444 Cl, 0.2 O So I would divide them all by Cl right and then round up to the nearest integer?

OpenStudy (anonymous):

so K = 1, O =I got .5 and I have no idea how to get Cl

OpenStudy (aaronq):

so K=1, then Cl=1, and O=2 \(KClO_2\)

OpenStudy (anonymous):

Hmm I think I messed up somewhere...

OpenStudy (anonymous):

That KClO2 formula is wrong...I don't think I used the 12.8 correctly

OpenStudy (aaronq):

so your masses were K 4.09 g Cl 3.71 g O = 12.82 g- (4.09 g + 3.71 g)

OpenStudy (aaronq):

what did you get for moles?

OpenStudy (anonymous):

I got the 0.2

OpenStudy (aaronq):

0.2 for moles of oxygen?

OpenStudy (anonymous):

Yeah, is that where I messed up?

OpenStudy (aaronq):

what mass did you use?

OpenStudy (anonymous):

7.8/15.999 = .4875. Is this the actual answer for moles?

OpenStudy (aaronq):

i don't think there are 7.8 grams of oxygen

OpenStudy (anonymous):

Oops forgot to subtract it from 12.82, so 5.2/15.999= .325

OpenStudy (aaronq):

now divide

OpenStudy (anonymous):

.10465444/.325 = .322

OpenStudy (aaronq):

no divide everything by the lowest number

OpenStudy (anonymous):

.10465444 is the lowest number

OpenStudy (aaronq):

you didn't divide by the lowest number here: .10465444/.325 = .322

OpenStudy (anonymous):

Ok I'm lost again

OpenStudy (aaronq):

okay, lets step back. what are your mole values?

OpenStudy (anonymous):

.10465444 for Cl, .10487179 for K, and .325 for O

OpenStudy (aaronq):

okay, now divide everything by the lowest number

OpenStudy (anonymous):

.10465444/.10487179 = .99 = 1, and .1046544/.325 = .322

OpenStudy (anonymous):

Now what?

OpenStudy (toxicsugar22):

Tokyo down can you hel. Me

OpenStudy (toxicsugar22):

In chem it is posted there

OpenStudy (aaronq):

you're still doing this wrong: ".1046544/.325 = .322" i asked you to divide by the smallest number, as in: .325/.1046544

OpenStudy (anonymous):

ok so 3.10545831 .1046/.1068 = 1.00207

OpenStudy (aaronq):

so your final formula is \(KClO_3\)